# ReCo-Diff: Explore Retinex-Based Condition Strategy in Diffusion Model for Low-Light Image Enhancement

### v1

@@ Abstract
Low-light image enhancement (LLIE) has achieved promising performance by employing conditional diffusion models. Despite the success of some conditional methods, previous methods may neglect the importance of an explicit formulation of task-specific condition strategy, resulting in suboptimal visual outcomes. In this study, we propose ReCo-Diff, a novel approach that incorporates Retinex-based prior as an additional pre-processing condition to regulate the generating capabilities of the diffusion model. ReCo-Diff first leverages a pre-trained decomposition network to produce initial reflectance and illumination maps of the low-light image. Then, an adjustment network is introduced to suppress the noise in the reflectance map and brighten the illumination map, thus forming the learned Retinex-based condition. The condition is integrated into a refinement network, implementing Retinex-based conditional modules that offer sufficient guidance at both feature- and image-levels. By treating Retinex theory as a condition, ReCo-Diff presents a unique perspective for establishing an LLIE-specific diffusion model. Extensive experiments validate the rationality and superiority of our ReCo-Diff approach. The code will be made publicly available.
1 Introduction
@@
Figure 1: Visual comparisons among recent DiffLL [14], PyDiff [57], and our ReCo-Diff on an low-light image from the LOL-v2 dataset. Previous diffusion-based methods exhibit detail loss and color distortion. Our method properly maintains color constancy and generates realistic textures thanks to the introduction of the superior Retinex-based condition strategy.
@@
Low-light photography is quite prevalent in the real world due to inherent environmental or technology restrictions, such as inadequate lighting and limited exposure times. Images captured in weakly illuminated conditions not only have low visibility for human perception but also are unsuitable for subsequent multimedia computing and downstream vision tasks, such as object detection [3, 19] and semantic segmentation [6]. Therefore, a large number of methods for low-light image enhancement (LLIE) are proposed to expose hidden features in low-light images and avoid performance degradation in subsequent vision tasks.
@@
Thanks to the development of diffusion models (DMs) [13, 33], numerous diffusion-based studies have been conducted for image restoration tasks [31, 28], with the goal of facilitating texture recovery. Previous diffusion-based image restoration methods choose to concatenate the low/high-quality images [31, 30] and extract features and priors [8, 10] as conditions. Furthermore, conditional DMs have already been introduced in LLIE tasks and thus several successful attempts have emerged [48, 14, 37, 57, 49]. It is worth noting that sampling efficiency is a common difficulty for the diffusion model, where limited sampling steps result in subpar generation fidelity. Therefore, Jiang et al[14] adopt 2D discrete wavelet transformations and utilize the low-resolution coefficients as conditions for faster inference speed. However, without the explicit modeling of color information, the results are unpleasing as shown in Fig. 1(a). To better adapt to LLIE task, the color map [49, 34, 57] and Retinex model [48] are introduced into the diffusion process. [34] and [49] try to maintain color consistency, while they only apply an invariant color map and gain limited ability of enhancement. Zhou et al. [57] propose a pyramid diffusion model (PyDiff) to perform the reverse process by progressively increasing resolution and adjusting the color after denoising through a global corrector, which improves the visual quality to a certain extent as shown in  Fig. 1(b). However, the visual results show limitations in preserving color and details, which may be solved by the solid Retinex theory. Thus, to facilitate the physical interpretability, Diff-Retinex [48] explores the possibility of establishing the Retinex-based diffusion model. The denoising network inputs reflectance and illumination maps as conditional images, and a consistent network is proposed to preserve the content, similar to PyDiff [57]. However, directly applying the decomposed maps as conditions is expected to be imperfect since Retinex theory is essentially an ideal model and the corruptions in reflectance and illumination need to be considered [1]. Although numerous studies have recognized the significance of auxiliary guidance in the diffusion process, they fail to explore a suitable strategy and still produce unfavorable visual results. Thus, the condition strategy has the potential for further improvement in establishing an LLIE-specific diffusion model with significant capability of enhancement.
@@
To address the aforementioned issues, we propose a novel Retinex-conditioned diffusion model, ReCo-Diff, specially designed for the LLIE task. From a new perspective, we combine physical models and the conditioning technique to produce better results. Specifically, our ReCo-Diff learns the Retinex-based prior as extra pre-processing conditions in the diffusion model and refines the enhanced output by integrating the Retinex-based conditions as guidance, instead of processing the reflectance and illumination maps directly [48]. In each step, our ReCo-Diff conducts Retinex-based condition learning and conditional image refinement successively. In the stage of condition learning, we first utilize a pre-trained network (DNet) to produce initial decomposed maps. The adjustment network (ANet) improves the initial maps by suppressing noise in reflectance and adjusting exposure in illumination, and the outputs serve as reliable Retinex-based conditions for better refinement. Then, different from simply using color maps as conditional inputs [49, 57, 34], our ReCo-Diff incorporates the Retinex-based conditions into the denoising U-Net (UNet) for better conditional guidance. Furthermore, unlike [57, 48] directly using stacked convolutional layers to improve the enhanced image, we propose a refinement network (RNet) to better preserve color and contents based on the Retinex theory, which separates the illumination while maintaining the color and details in the reflectance. We reformulate the Retinex model into a residual manner to alleviate the corruptions and thus design the feature- and image-level Retinex-conditioned modules (F/IRCM) in the RNet. The physically inspired conditions and the modified Retinex model for refinement endow our ReCo-Diff with the superior capability of enhancement. Finally, as shown in Fig. 1(c), our ReCo-Diff provides the most pleasing result by using the novel task-specific condition strategy.
@@
  • We propose a novel diffusion-based method for LLIE with a Retinex-based condition strategy, exploring the role of physical models in controlling the generation capabilities of diffusion models.
  • We formulate our ReCo-Diff from two aspects: the condition learning stage that improves initial decomposed maps and produces reliable Retinex-based conditions; and the conditional refinement stage that exploits the Retinex-based conditions to guide the diffusion process for better preserving the color and contents.
@@
  • Extensive experiments on representative benchmarks demonstrate that the Retinex-based condition strategy and the well-designed learning and refinement process lead to the superior performance of our ReCo-Diff.
@@
Figure 2: Overview of our Retinex-based conditional diffusion model (ReCo-Diff). Each step contains two stages: (a) In the learning stage, the pre-trained DNet provides the initial decomposed maps and the ANet adjusts and outputs reliable Retinex-based conditions. (b) In the refinement stage, the learned conditions are incorporated into UNet as extra inputs and are incorporated into RNet through F/IRCMs to preserve the color and content. FRCM and IRCM respectively receive the features and adjusted maps from ANet, detailed in Fig. 3.
2 Related Work
Low-light Image Enhancement. Numerous works introduce the Retinex theory into deep neural networks [39, 52, 53, 40, 1, 48, 5, 7, 21, 22]. The Retinex-Net [39] is the most inspiring method combining physical model and DNNs. Then, Zhang et al. proposes KinD [52] and KinD++ [53] to provide more effective solutions. Rather than complex multi-stage training pipeline, Fu et al[5] and Cai et al[1] explore the possibility of end-to-end frameworks and achieve significant performance improvement.
Without Retinex theory, recent works concentrate on directly end-to-end manners [42, 26, 55, 36, 50, 43, 29, 44, 16]. LLNet [27] inspires the emergence of end-to-end methods. Primarily, supervised methods show promising capability of enhancement. For reducing color deviation,  [45, 20, 54, 41] adopt 3DLUT and histogram to preserve color consistency. In [43, 44], the SNR-aware prior and the structure-aware features are taken as guidance to produce realistic results. Recently, the Ultra-High-Definition (UHD) images become popular. LLformer [35] and UHDFour [24] are proposed to enhance UHD images and release UHD datasets to promote the following research. Furthermore, semi-supervised [46, 29], unsupervised [15, 7] and zero-shot learning [9, 56, 23] are valuable when training images are limited.
@@ 2 Related Work
Diffusion-Based Image Restoration and Low-Light Image Enhancement. Diffusion models [13, 33] show promising capability in image generation tasks [32, 2]. For solving image processing tasks, diffusion-based methods employ conditional mechanism to incorporate the distorted images as guidance [25], such as colorization [31], super-resolution [32], restoration [28, 18, 4], and LLIE [34, 48, 49, 37, 57, 14]. To realize efficient diffusion-based LLIE, DiffLL [14] uses wavelet transformation to decrease the input size and a high-frequency restoration module to maintain the details. PyDiff [57] directly down-samples the image in early steps to speed up the sampling process. Considering the characteristics of LLIE task, LLDiffusion [34] and CLE-Diffusion [49] use the color map as extra conditional input to preserve the color information. Furthermore, PyDiff [57] proposes a global corrector to alleviate color degradation. By introducing Retinex theory, Diff-Retinex [48] decomposes the image and utilizes the reflectance and illumination maps as conditional images to form the guidance.
@@
Although existing diffusion-based LLIE methods obtain good performance, they are limited by insufficient conditional guidance and produce unsatisfactory color and details. In this paper, we investigate a new condition strategy and propose ReCo-Diff to control the generative diffusion model by forming a Retinex-based conditional process, exploring a novel perspective for diffusion-based LLIE and achieving promising performance.
3 Method
@@
The overview is shown in Fig. 2. We present the Retinex-based condition strategy to explore an effective diffusion-based method for LLIE. We introduce the conditional DDPM in Sec. 3.1 and present our ReCo-Diff in Sec. 3.2.
3.1 Conditional Denoising Diffusion Model
@@
To deal with image restoration tasks, condition strategies are proposed to develop the conditional DDPM [25]. The conditional DDPM also generates a target image $\mathbf{x}_{0}$ from a pure noise image $\mathbf{x}_{T}$ and refines the image through successive iterations. Unlike DDPM, the low-quality images are utilized as conditional inputs in conditional DDPM [32]. Thus, the conditional inference is defined as a reverse Markovian process:
@@
$\begin{split}p_{\theta}(\mathbf{x}_{0:T}|\mathbf{y})&=p(\mathbf{x}_{T})\prod\nolimits_{t=1}^{T}p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{y}),\\ p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{y})&=\mathcal{N}(\mathbf{x}_{t-1}\mid\mu_{\theta}({\mathbf{x}}_{t},\mathbf{y},\gamma_{t}),\sigma_{t}^{2}\mathbf{I}),\end{split}$ (1)
@@
where $\gamma_{t}=\prod_{i=1}^{t}\alpha_{i}$ and $\alpha_{1:T}$ are scale parameters. Due to the absence of high-quality image $\mathbf{x}_{0}$ in inference, ${\bm{\epsilon}}_{\theta}$ estimates noise and approximates $\mathbf{x}_{0}$ as:
@@
$\hat{\mathbf{x}}_{0}=\frac{1}{\sqrt{\gamma_{t}}}\left(\mathbf{x}_{t}-\sqrt{1-\gamma_{t}}\,\bm{\epsilon}_{\theta}(\mathbf{x}_{t},\mathbf{y},\gamma_{t})\right).$ (2)
@@
Therefore, the mean of $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{y})$ can be parameterized by applying the $\hat{\mathbf{x}}_{0}$ into the posterior distribution [13]:
@@
$\!\mu_{\theta}(\mathbf{y},{\mathbf{x}}_{t},\gamma_{t})=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\gamma_{t}}}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},\mathbf{y},\gamma_{t})\right).$ (3)
@@
$\mathbb{E}_{\mathbf{x}_{0},\mathbf{y},t,\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\lVert\bm{\epsilon}_{t}-\bm{\epsilon}_{\theta}(\mathbf{x}_{t},\mathbf{y},\gamma_{t})\rVert^{2}_{2}].\vspace{-0.05cm}$ (4)
3.2 Retinex-Based Conditional Diffusion Model

### v2

# JoReS-Diff: Joint Retinex and Semantic Priors in Diffusion Model for Low-light Image Enhancement

@@ Abstract.
Low-light image enhancement (LLIE) has achieved promising performance by employing conditional diffusion models. Despite the success of some conditional methods, previous methods may neglect the importance of a sufficient formulation of task-specific condition strategy, resulting in suboptimal visual outcomes. In this study, we propose JoReS-Diff, a novel approach that incorporates Retinex- and semantic-based priors as the additional pre-processing condition to regulate the generating capabilities of the diffusion model. We first leverage pre-trained decomposition network to generate the Retinex prior, which is updated with better quality by an adjustment network and integrated into a refinement network to implement Retinex-based conditional generation at both feature- and image-levels. Moreover, the semantic prior is extracted from the input image with an off-the-shelf semantic segmentation model and incorporated through semantic attention layers. By treating Retinex- and semantic-based priors as the condition, JoReS-Diff presents a unique perspective for establishing an diffusion model for LLIE and similar image enhancement tasks. Extensive experiments validate the rationality and superiority of our approach.
low-light image enhancement, conditional diffusion model, Retinex model, semantic guidance
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3680876isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Image processing
1. Introduction
@@
Figure 1. Visual comparisons among recent DiffLL (Jiang et al., 2023), PyDiff (Zhou et al., 2023), and our JoReS-Diff on LOL-v2 dataset. Previous diffusion-based methods exhibit detail loss and color distortion. Our method properly maintains color constancy and generates realistic textures thanks to the introduction of the superior Retinex and semantic priors.
@@
Low-light photography is quite prevalent in the real world due to inherent environmental or technology restrictions. Low-light images are not only harmful for human perception but also for downstream vision tasks, such as object detection (Deng et al., 2019; Kennerley et al., 2023) and semantic segmentation (Fu et al., 2019). Thus, various methods for low-light image enhancement (LLIE) are proposed to improve the quality of low-light images.
@@
Thanks to the development of diffusion models (DMs) (Ho et al., 2020; Song et al., 2020), numerous diffusion-based studies have been conducted for image restoration tasks (Saharia et al., 2022a; Luo et al., 2023), with the goal of facilitating texture recovery. Previous diffusion-based image restoration methods choose to concatenate the low/high-quality images (Saharia et al., 2022a; Niu et al., 2023) and extract features and priors (Gao et al., 2023; Guo et al., 2023b) as conditions. Furthermore, conditional DMs have already been introduced in LLIE tasks and thus several successful attempts have emerged (Yi et al., 2023; Jiang et al., 2023; Wang et al., 2023a; Zhou et al., 2023; Yin et al., 2023). It is worth noting that sampling efficiency is a common difficulty for the diffusion model. Therefore, Jiang et al. .  (Jiang et al., 2023) adopt 2D discrete wavelet transformations and utilize the low-resolution coefficients as conditions for faster inference speed. However, without the explicit modeling of color information, the results are unpleasing as shown in Fig. 1 (a). To better adapt to LLIE task, the color map (Yin et al., 2023; Wang et al., 2023b; Zhou et al., 2023) and Retinex model (Yi et al., 2023) are introduced into the diffusion process. (Wang et al., 2023b) and (Yin et al., 2023) try to maintain color consistency, while they only apply an invariant color map and gain limited ability of enhancement. PyDiff (Zhou et al., 2023) proposes a pyramid resolution setting to perform faster reverse process and adjusts the color through a global corrector, which improves the visual quality to a certain extent as shown in  Fig. 1(b). However, the visual results show limitations in preserving color and details. Thus, Diff-Retinex (Yi et al., 2023) explores the possibility of establishing the Retinex-based diffusion model and produces better color and content. The denoising networks input reflectance and illumination maps as conditional images, and consistent networks are proposed to preserve the content . However, directly applying the decomposed maps as conditions is expected to be imperfect since Retinex theory is essentially an ideal model and the corruptions in reflectance and illumination need to be considered (Cai et al., 2023). Although numerous studies have recognized the significance of auxiliary guidance in the diffusion process, they fail to explore a suitable strategy and still produce unfavorable visual results. Thus, the condition strategy has the potential for further improvement in establishing an LLIE-specific diffusion model with significant capability of enhancement.
Figure 2. Comparison between Diff-Retinex (Yi et al., 2023) and ours. Diff-Retinex still relies on the original decomposition and multiplication process uses two diffusion models to process the decomposed maps, which are multiplied as output. Our method is end-to-end and uses both Retinex and semantic priors, which are integrated into one single diffusion model. We also propose RNet to fully exploit the Retinex prior.
@@
To address the aforementioned issues, we propose a novel Retinex- and semantic-conditioned diffusion model, JoReS-Diff, specially designed for the LLIE and similar image enhancement tasks. From a new perspective, we combine physical models and semantic guidance as joint priors for the manipulation of the diffusion process. Retinex prior: JoReS-Diff incorporates the Retinex priors as extra pre-processing condition in the diffusion model, which is different from Diff-Retinex (Yi et al., 2023) as shown in Fig. 2. Specifically, JoReS-Diff conducts Retinex-based condition learning and conditional image refinement stages. In the former stage, we first utilize a pre-trained network (DNet) to produce initial decomposed maps. The adjustment network (ANet) suppresses noise in reflectance and adjusts exposure in illumination and outputs more reliable Retinex-based condition for better refinement. Then, JoReS-Diff incorporates the decomposed maps into the denoising U-Net (UNet) through Retinex attention layers for better conditional guidance. Furthermore, unlike (Zhou et al., 2023; Yi et al., 2023) directly adding a network after denoising process to improve image quality, we propose a refinement network (RNet) to better preserve color and contents based on the Retinex theory, which separates the illumination while maintaining the color and details in the reflectance. We reformulate the Retinex model into a residual manner to alleviate the corruptions and thus design the feature- and image-level Retinex-conditioned modules (F/IRCM) in the RNet. Semantic prior: Although the Retinex prior provides color and detail refinement, the diffusion model still suffers from unnatural textures (Zhang et al., 2023) without considering semantic guidance. Thus, we incorporate the semantic prior for better controlling the generation ability of the diffusion model. To be specific, we extract the semantic prior from the input with an off-the-shelf semantic segmentation model and incorporate it through semantic attention layers like (Wu et al., 2023). We both use self-attention and cross-attention for inherent structure preservation and semantic consistency simultaneously (Liu et al., 2024). Finally, as shown in Fig. 1(c), our JoReS-Diff provides the most pleasing result by using the novel joint condition strategy.
@@
  • We propose a novel diffusion-based method for image enhancement with joint Retinex and semantic priors, exploring the role of physical models and semantic guidance in controlling the generation capabilities of diffusion models.
  • We propose condition learning and conditional refinement stages to integrate Retinex prior and preserve color and content consistency. We also introduce semantic prior by semantic attention layers to control the generation ability of diffusion model and reserve semantic consistency.
@@
  • Extensive experiments on representative benchmarks demonstrate that the joint Retinex and semantic priors and the well-designed interaction mechanism lead to the superior performance of our JoReS-Diff.
@@
Figure 3. Overview of our Retinex- and semantic-based conditional diffusion model (JoReS-Diff). (a) The introduction of Retinex prior contains two stages. In the learning stage, DNet provides the initial decomposed maps and ANet outputs reliable Retinex-based conditions; In the refinement stage, the conditions ${R_{t}}\!^{\prime},{L_{t}}\!^{\prime},F_{t}$ are incorporated into UNet and RNet through Retinex attention layers and F/IRCMs (detailed in Fig. 4) to preserve the color and content. (b) The semantic prior $\mathbf{c}_{seg}$ is extracted by a pre-trained segmentation model and then integrated into UNet through semantic attention layers.
2. Related Work
@@
Diffusion models (Ho et al., 2020; Song et al., 2020) show promising capability in image generation tasks (Saharia et al., 2022b; Choi et al., 2021). For solving image processing tasks, diffusion-based methods employ conditional mechanism to incorporate the distorted images as guidance (Li et al., 2023b), such as colorization (Saharia et al., 2022a), super-resolution (Saharia et al., 2022b), restoration (Luo et al., 2023; Kawar et al., 2022; Fei et al., 2023), and LLIE (Wang et al., 2023b; Yi et al., 2023; Yin et al., 2023; Wang et al., 2023a; Zhou et al., 2023; Jiang et al., 2023). To realize efficient diffusion-based LLIE, DiffLL (Jiang et al., 2023) uses wavelet transformation to decrease the input size and a high-frequency restoration module to maintain the details. PyDiff (Zhou et al., 2023) directly down-samples the image in early steps to speed up the sampling process. Considering the characteristics of LLIE task, LLDiffusion (Wang et al., 2023b) and CLE-Diffusion (Yin et al., 2023) use the color map as extra conditional input to preserve the color information. Furthermore, PyDiff (Zhou et al., 2023) proposes a global corrector to alleviate color degradation. By introducing Retinex theory, Diff-Retinex (Yi et al., 2023) decomposes the image and utilizes the reflectance and illumination maps as conditional images to form the guidance.
@@
Although existing diffusion-based LLIE methods obtain good performance, they are limited by insufficient conditional guidance and produce unsatisfactory color and details. In this paper, we investigate a new condition strategy and propose JoReS-Diff to control the generative diffusion model by forming a Retinex- and semantic-based conditional process, exploring a novel perspective for diffusion-based LLIE and similar image enhancement tasks.
3. Method
@@
The overview is shown in Fig. 3. We present the Retinex- and semantic-based condition strategy to explore an effective diffusion-based method for LLIE. We introduce the conditional DDPM in Section 3.1 and present our JoReS-Diff in Section 3.2 and  Section 3.3.
3.1. Conditional Denoising Diffusion Model
@@
To deal with image restoration tasks, condition strategies are proposed to develop the conditional DDPM (Li et al., 2023b). The conditional DDPM also generates a target image $\mathbf{x}_{0}$ from a pure noise image $\mathbf{x}_{T}$ and refines the image through successive iterations. Unlike DDPM, the low-quality images are utilized as conditional inputs in conditional DDPM (Saharia et al., 2022b). Thus, the conditional inference is defined as a reverse Markovian process:
@@
(1) $\begin{split}p_{\theta}(\mathbf{x}_{0:T}|\mathbf{y})&=p(\mathbf{x}_{T})\prod\nolimits_{t=1}^{T}p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{y}),\\ p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{y})&=\mathcal{N}(\mathbf{x}_{t-1}\mid\mu_{\theta}({\mathbf{x}}_{t},\mathbf{y},\gamma_{t}),\sigma_{t}^{2}\mathbf{I}),\vspace{-0.25cm}\end{split}$
@@
where $\gamma_{t}=\prod_{i=1}^{t}\alpha_{i}$ and $\alpha_{1:T}$ are scale parameters, $\mathbf{y}$ denotes the low-light input. Due to the absence of high-quality image $\mathbf{x}_{0}$ in inference, ${\bm{\epsilon}}_{\theta}$ estimates noise and approximates $\mathbf{x}_{0}$ as:
@@
(2) $\hat{\mathbf{x}}_{0}=\frac{1}{\sqrt{\gamma_{t}}}\left(\mathbf{x}_{t}-\sqrt{1-\gamma_{t}}\,\bm{\epsilon}_{\theta}(\mathbf{x}_{t},\mathbf{y},\gamma_{t})\right).\vspace{-0.15cm}$
@@
Therefore, the mean of $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{y})$ can be parameterized by applying the $\hat{\mathbf{x}}_{0}$ into the posterior distribution (Ho et al., 2020):
@@
(3) $\!\mu_{\theta}(\mathbf{y},{\mathbf{x}}_{t},\gamma_{t})=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\gamma_{t}}}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},\mathbf{y},\gamma_{t})\right).\vspace{-0.15cm}$
@@
(4) $\mathbb{E}_{\mathbf{x}_{0},\mathbf{y},t,\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\lVert\bm{\epsilon}_{t}-\bm{\epsilon}_{\theta}(\mathbf{x}_{t},\mathbf{y},\gamma_{t})\rVert^{2}_{2}].\vspace{-0.15cm}$
3.2. Retinex Prior Incorporation